### v7

# Forecasting Treatment Response with Deep Pharmacokinetic Encoders

Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, forecasting is challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties for each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data. Our hybrid global-local architecture improves over patient-specific models by 15.8%  on average. Additionally, our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
keywords:
Multi-Input Forecasting, Healthcare, Pharmacokinetics, Deep Learning
1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the medication dose is usually known, one cannot directly observe the plasma concentration of the drug, as illustrated in Fig. 1. Consequently, the impact of dynamic changes in treatment resulting from drug absorption and elimination remains hidden.
Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and varies over time.
We consider the blood glucose forecasting task while taking into account latent, time-dependent treatment effects. Patients with diabetes must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions for hyperglycemia indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospital admissions suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year. Capturing the dynamic effects of insulin is crucial for accurately predicting its impact on blood glucose levels to anticipate and proactively address abnormal levels to prevent adverse patient events.
The main contributions of this paper are:
  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
  • Pharmacokinetic (PK) Encoder. We introduce a novel deep learning model-agnostic module that generates plasma concentration profiles to capture time-dependent treatment effects.
  • Multiple Dose Effects Framework. We provide a computationally efficient framework for encoding the effects of multiple doses. We verify this approach mathematically under the assumption of linear PK.
  • State-of-the-Art Results. We obtain significant accuracy improvements on large scale simulated and real-world patient blood glucose forecasting datasets.
1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement within the body. For example, PK provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The degree of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
2 Related Work
Blood Glucose Forecasting. Neural Forecasting has previously been applied to the field of blood glucose forecasting, outperforming the accuracy of classic statistical methods. RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been utilized in proposed models and as baselines.
Blood Glucose Treatment Effect Modeling. Previous studies have successfully modeled treatment effects on patient blood glucose levels. Miller et al. present one of the first hybrid statistical-and-physiological model that combines glucose-insulin dynamics and deep learning (Miller et al., 2020). They present the Deep T1D Simulator (DTD-Sim) model, which uses a deep state-space model to describe parameters representing insulin and carbohydrate absorption rates and sensitivities of the UVA/Padova simulator that simulates blood glucose levels. Odnoblyudova et al. demonstrated enhanced accuracy in predicting blood glucose levels using Multi-output Gaussian Processes (MOGPs) to model composite effects of carbohydrate and fat components from meals. (Odnoblyudova et al., 2023).
Despite recent advancements in forecasting treatment effects, significant challenges remain, including:
Sparse data.
Limited treatment data can hinder deep learning models from accurately learning treatment responses (Rubin-Falcone et al., 2022; Hussain et al., 2021). To address this challenge, we integrate domain knowledge into the deep learning models, allowing them to more effectively encode the time-dependent effects of these variables and improve forecasting performance.
Composite treatment effects.
Unlike meals that consist of various components such as carbohydrates and fats, insulin is a homogenous protein hormone composed of a single type of molecule. Therefore, modeling its component effects is not applicable (Odnoblyudova et al., 2023). Instead, we aim to model patient-specific holistic effects of the drug as well as composite effects of multiple doses.
Inflexible models.
Previous glucose forecasting models, like the UVA/Padova simulator, are specific to blood glucose dynamics and inflexible to other forecasting targets and variables (Visentin et al., 2016; Miller et al., 2020). Our approach leverages a deep learning model to generate predictions. This distinction provides greater flexibility, as our forecasting model is not constrained by a specific physiological model and can accommodate various deep learning models and forecasting targets and inputs.
Model application scope.
Prior work that integrates PK domain knowledge in deep learning models has been used in the context of predicting disease progression for cancer and COVID-19 with long-term horizons of interest (Hussain et al., 2021; Qian et al., 2021). In contrast, we leverage linear PK assumptions to create a framework that can be directly applied to various medications used in clinical practice and is easily adapted to other treatments.
Individualized networks.
Studies often use individualized networks, limiting cross-learning. Our novel global-local architecture allows efficient training of a more accurate global model applicable to multiple patients.
3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with GPU memory constraints. the duration of the rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (ins, 2018, ). Therefore, the selection of a 10-hour lag time can fully cover multiple dose durations, allowing the model to learn treatment effects across various dosages and durations.
3.1 PK Encoder
Medication administrations are represented as instantaneous, sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate that specific to patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from pharmacokinetics literature as outlined in Section 1.1 and Appendix B.
Figure 3: Medication administrations are represented as sparse variables in time series data. We propose a pharmacokinetic (PK) encoder to effectively capture time-dependent plasma drug concentration (Conc.).
Under linear PK assumptions, PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose, x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude but does not alter the PK parameters or shape of the curve.
Let $N$ be the number of subjects in the data. The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training. The initial k values are tuned as model hyperparameters.
The encoder is trained end-to-end with the model. During model training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder than replaces sparse medication features in the training window with the corresponding concentration curve. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded in the time prior to $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include plasma treatment effects of CHO because, similar to drugs, it undergoes absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
The PK encoder training framework is outlined in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables alongside modeling the target variable.
3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 11. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters. Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$. PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs. During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
3.3 Modeling Dose-Dependent Effects
A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017). A characteristic of this assumption is that PK parameters are constant and do not change with dose. In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site. Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose. For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018). This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
CGMs administer the same type of insulin (rapid-acting) for basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 9 and 10 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity. By treating each dose range as a separate entity, we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity of a full nonlinear model while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
Treating each dose as a different entity is an approximation and may not capture all nuances of dose-dependent kinetics. All physiological models are approximations of complex biological processes. Modeling dose-dependent effects in a more granular manner could potentially provide more precise results if such effects are present. However, our aim is to develop minimal models validated under established linear PK assumptions from the pharmacology literature, striking a balance between simplicity and accuracy. Our framework ensures broader applicability across various medications where a majority of drugs are assumed to adhere to linear PK (Rosenbaum, 2017).
The equations presented in the paper are shown in a patient-specific manner to highlight that the PK encoder generates individualized information. However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.
3.4 Models
To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines. We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model. Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al. (2020); Challu et al. (2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al. (2014)), Long-Short Term Memory (LSTM; Sak et al. (2014)), and Temporal Convolution Network with an MLP Decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)). Implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022) and StatsForecast (Garza et al., 2022) libraries. More information on these models and related prior work that leverages these are provided in Appendix A.2. Details on model training and hyperparameter selection are provided in Appendix A.3
3.5 Baselines
We compare our hybrid global-local architecture with the PK encoder to both local and global baselines:
Global Baseline Models.
Models were trained on all patient data to create a single global model:
  • Baseline. Models are trained with the univariate blood glucose signal to demonstrate the forecasting improvement from using exogenous variables, supporting our PK encoder.
  • Baseline Exogenous (Exog.). Models are trained with the univariate blood glucose signals, static features, and sparse exogenous features (CHO and insulin dose values).
  • Baseline Exogenous (Exog.) - ST. Models are trained with the univariate blood glucose signal, static features, and sparse exogenous features transformed via the sum total (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
Local Baseline Exogenous Models.
Patient-specific baseline exogenous models were compared against our hybrid global-local PK encoder model.
Local PK Encoder Models.
Patient-specific models using the PK encoder were compared against our hybrid global-local PK encoder model.
3.6 Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models. Formulations for these metrics are provided in Appendix A.4. We evaluate model performance across all points within the specified time horizon. We also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
4 Cohort
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. Blood glucose time series plots for 6 subjects are shown Fig. 7 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 9 in Appendix A.1.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The data includes basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 10 in Appendix A.1.
More dataset information is in Appendix A.1.
5 Results
5.1 Results on Simulated Data
For the simulated dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous model for forecasts evaluated at all values (t-test p-value: 6.6e-6) as shown in as shown in Table 1. For evaluations across all values, the NHITS-PK model provides 3.5% improvement, on average, and up to 9.2% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK model provides 5.2% improvement, on average, and up to 16.4% improvement for individual patients over the baseline NHITS sparse exogenous model. Additionally, the NHITS-PK model outperforms the NHITS-ST model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 6 in Appendix C.5.
5.2 Results on Real Data
Our hybrid global-local architecture produces models that significantly outperform local models by 15.8%, on average (t-test p-value: 1.7e-5), and up to 31.3% improvement for individual OhioT1DM patients. We show that the NHITS-PK hybrid global-local model outperforms NHITS-PK models developed for individual OhioT1DM patients in Fig. 12. Moreover, this improvement in forecast accuracy is consistent across all patients as shown in Fig. 5 and in Table 4 in Appendix C.1. Additionally, the NHITS-PK model with the hybrid global-local architecture also improves MAE by 14.8%, on average, for patients over the patient-specific NHITS sparse exogenous baseline models (t-test p-value: 1.7e-6).
For the OhioT1DM dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous baseline model for forecasts evaluated at all values (t-test p-value: 5.0e-3) and critical values (t-test p-value: 3.5e-3) as shown in Table 1. For evaluations across all values, The NHITS-PK model provides 1.3%  on average, and up to 3.5% improvement for individual patients over the baseline NHITS sparse exogenous model. For evaluations across critical values, the NHITS-PK provides 1.6%  on average, and up to 4.9% improvement for individual patients over the baseline NHITS sparse exogenous model. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 7 in Appendix C.5. Fig. 13 in Appendix C.2 shows model forecasts during hyperglycemic events for four randomly selected OhioT1DM patients.
Figure 5: The NHITS-PK hybrid global-local model (blue, solid) provides more accurate predictions than local NHITS-PK models (purple, dashed).
Figure 6: In addition to reducing forecasting error during hyperglycemic events, our NHITS-PK encoder enables improved lead time predictions of hypertensive events (blood glucose $\geq$ 180) for 9 of 12 OhioT1DM patients as measured over the 8 trials.
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
Simulated Dataset
Model All Values Critical Values
Baseline Exog. Baseline Exog.
ETS 9.450 13.457
MLP 10.589 8.433 14.699 10.600
RNN 13.703 20.190 21.433 35.955
LSTM 13.557 10.321 20.537 13.550
TCN 12.096 10.109 17.709 12.660
NHITS 8.268 7.304 11.114 9.094
NHITS-ST 7.055 8.648
NHITS-PK 7.037 8.492
OhioT1DM Dataset
ETS 9.022 10.028
MLP 11.065 11.120 12.630 12.698
RNN 11.325 11.795 13.627 13.552
LSTM 9.811 9.467 11.252 11.178
TCN 11.525 10.760 13.647 12.678
NHITS 9.060 8.873 10.337 10.137
NHITS-ST 8.958 10.131
NHITS-PK 8.758 9.965
6 Discussion
Hybrid Global-Local Architecture We demonstrate the efficacy of our novel hybrid global-local model, which produces forecasts with significantly smaller errors than patient-specific models across all patients. Our hybrid architecture not only improves accuracy but also enables global models to learn relevant patient-specific information, outperforming global model baselines.
A hybrid global-local model offers practical benefits beyond improved forecasting accuracy as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model can also be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
PK Encoder As a model-agnostic extension, our novel PK encoder significantly reduces forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts such events earlier, up to 3.6 minutes when compared to the NHITS baseline model, for individual OhioT1DM patients as shown in Fig. 6. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
Our results also highlight the importance of using domain knowledge to incorporate exogenous variables into models. Models with sparse features may underutilize valuable information, and feature-engineered representations not informed by domain knowledge, like the ‘sum total’ approach, can underperform compared to baselines. Moreover, Appendix C.4 shows that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset. As such, dose-dependent models, such as proposed in section 3.3, may be needed for more realistic data simulations and models.
Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing or drug selection. It can also be adapted for other medication administration types, such as oral or intravenous injections.
Limitations. We equate $\sigma$ to our k parameter in the log-normal function and hold $\mu$ constant at 1. This simplification reduces the number of potential concentration curve shapes the model can learn. Tuning both $\sigma$ and $\mu$ could further improve forecasts by enabling more flexible concentration curves.
Future Work. Future work will explore fine-tuning global forecasting models with patient-specific parameters for new patients with limited historical data as well as modeling dose-dependent effects using more dynamic nonlinear PK parameter frameworks. Future work will include applying our approach to additional patient cohorts, such as those with T2DM, to further validate our findings.
[…]
This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
Appendix A Supplemental Methods
Data and Code Availability
We use open-source simulated and real-world data that is available to other researchers. The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository [Xie, 2018, Visentin et al., 2016]. The real-world data includes the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu [2020].
Implementations of models used in our work were obtained from the open-source Neuralforecast [Olivares et al., 2022] and StatsForecast [Garza et al., 2022] libraries. We implemented our method using a copy of the Neuralforecast repository to leverage their model training framework. Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
A.1 Data
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) [Xie, 2018, Visentin et al., 2016] provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to [Rubin-Falcone et al., 2022], the meal schedule used to generate simulated data was based on the Harrison-Benedict equation [Arthur Harris and Gano Benedict, 21919] with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
The ‘Dexcom’ continuous glucose monitor (CGM) option was used to obtain glucose readings at 5-minute intervals. The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio. However, we adapt the controller to deliver basal doses once per hour, consistent with the function of real insulin pumps used in the OhioT1DM dataset. The resulting simulated data consists of CGM (in mg/dL) values, insulin (units), and carbohydrates (in grams), referred to as CHO. We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses. Blood glucose time series for 6 simulated patients are shown in Fig. 7 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 9 in Appendix A.1.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020]. Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
Sparse exogenous features in the data include basal rate (in units per hour), temporary basal rate (units per hour), and bolus insulin (units), and CHO values. Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period. ‘Square dual’ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps. If multiple bolus administrations occurred within the same minute, the insulin units were summed. For Medtronic insulin pumps 530G and 630G versions, basal insulin is administered hourly based on the basal rate. For timestamps with overlapping basal rates, the last basal rate value was considered. When specified, the temporary basal rate superseded the basal rate.
We use the specified train and test partitions provided in the dataset. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Missing values were forward filled to prevent data leakage. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 10 in Appendix A.1.
Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 9 and 10 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 2: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
ID Original Preprocessed
Train Test Test
540 13109 3066 2691
(8.87%) (6.43%) (6.91%)
544 12671 3137 2691
(16.16%) (13.42%) (14.08%)
552 11097 3950 2691
(18.80%) (40.15%) (53.36%)
559 12081 2876 2691
(10.64%) (12.59%) (13.45%)
563 13098 2691 2691
(7.44%) (4.50%) (4.50%)
567 13536 2871 2691
(19.78%) (16.79%) (16.91%)
570 11611 2880 2691
(5.42%) (4.69%) (4.91%)
575 13103 2719 269
(9.44%) (4.74%) (4.79%)
584 13248 2995 2691
(8.29%) (11.02%) (11.93%)
588 13105 2881 2691
(3.55%) (3.12%) (3.34%)
591 12755 2847 2691
(14.96%) (3.06%) (3.23%)
596 13630 3003 2691
(20.20%) (8.66%) (9.66%)
A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies [Qian et al., 2021, Li et al., 2019a, Fox et al., 2018, Rabby et al., 2021, Rubin-Falcone et al., 2020, 2022, Miller et al., 2020]. Similarly, CNNs [Xie and Wang, 2020, Li et al., 2019a, b] and MLPs [McShinsky and Marshall, 2020, Miller et al., 2020] have been utilized in proposed models and as baselines. NHITS achieves state-of-the-art performance, surpassing RNN-based, transformer-based, and other deep learning time series decomposition models (such as N-BEATS [Oreshkin et al., 2020]) in long-horizon forecasting tasks [Challu et al., 2022].
Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. [Hyndman, 2008].
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation [Nair and Hinton, 2010, Fukushima, 1975, Rosenblatt, 1958].
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation [Challu et al., 2022].
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.[Elman, 1990, Cho et al., 2014]
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.[Sak et al., 2014]
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions [Bai et al., 2018, van den Oord et al., 2016].
A.3 Model Training and Hyperparameters
Models were trained using Adam optimizer [Kingma and Ba, 2017]. Optimal hyperparameters were selected based on a validation set with cross-validation. Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (8)
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
Hyperparameter Considered Values
Learning rate Loguniform(1e-4, 1e-1)
Batch size 4
Windows batch size’ 256
Dropout 0.0
Training steps 2000
Validation check steps 100
Early stop 5
patience steps
Random seed DiscreteRange(1, 10)
Initial $\textbf{k}_{\text{basal}}$1.0, 1.1, 1.2}
Initial $\textbf{k}_{\text{bolus}}$1.7, 1.8, 1.9}
Initial $\textbf{k}_{\text{CHO}}$1.7, 1.8, 1.9}
Table 3: Common hyperparameter search space
We tune additional specific hyperparameters for each architecture. For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the NHITS we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Metrics
We use the following metrics to evaluate model performance. Here, $H$ refers to the forecast horizon.
Mean Absolute Error (MAE)
$MAE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\frac{1}{H}\sum_{i=t+1}^{t+H}|y_{i}-\hat{y}_{i}|$
Root Mean Square Error (RMSE)
$RMSE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\sqrt{\frac{1}{H}\sum_{i=t+1}^{t+H}(y_{i}-\hat{y}_{i})^{2}}$
A.5 Propositions
Proposition A.1.
Under linear pharmacokinetic (PK) assumptions, the area under the concentration curve is directly proportional to the dose of the drug. As such, PK parameters, such as bioavailability, are assumed to be constant for a given drug. The bioavailability, $F$, can be assessed from the dose-normalized area under the concentration curve (AUC). When bioavailability is constant, we show that the AUC is bounded by the dose, $x$, given that $0\leq F\leq 1$,
$F$ $\propto\frac{\text{AUC}}{x}$ (9)
$n\cdot F$ $=n\cdot\frac{\text{AUC}}{x}$ (10)
AUC $=F\cdot x$ (11)
$\leq x$ (12)
where $n$ where is a constant of proportionality.
Proposition A.2.
Given some mathematical model, $C(t,\cdot)_{x}$, of concentration for a drug dose, $x$, that is a function of time, $t$, and pharmacokinetic parameters, $\cdot$, we show that the overall effect (concentration) of $n$ drug doses of the same drug is equal to the sum of the individual effects under linear pharmacokinetic assumptions:
[…]
$=\int_{0}^{T}\sum_{i=1}^{n}C(t,\cdot)_{x}dt$ (17)
Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
A.6 Algorithm
We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles. This framework can be implemented in any deep learning architecture. It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library [Liaw et al., 2018]. For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
Algorithm 1 PK Encoder Training.
$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
Result: k
$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
for each step i from 1 to E do
       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
for batch $b\in\mathcal{B}_{train}$ do
               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
       end for
      if i % validation step == 0 then
               $\ell(\textbf{k})\leftarrow 0$
for batch $b\in\mathcal{B}_{val}$ do
                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
             end for
            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
       end if
end for
return k
Figure 7: Time series for the first 10 days of 6 subjects from the simulated dataset.
Figure 8: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
Figure 9: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
Figure 10: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
Appendix B Pharmacokinetic Background
Concentration-time Profiles
From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects [Rosenbaum, 2017, Angelika et al., 2005]. Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile. Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves. The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig. 2, along with various patient-specific factors.
Insulin Analogs
Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action. Fig. 2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs [Angelika et al., 2005]. Generally, bolus doses are quick-acting and, thus, often taken before meals. Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day. However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin. For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes. Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
Bioavailability
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent [Søeborg et al., 2012]. However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
First-order Kinetics
First-order kinetics, a common assumption in pharmacokinetics, describes the rate at which a drug is eliminated from the body as being proportional to its concentration. This means that a constant fraction of the drug is eliminated per unit of time, regardless of its concentration levels. Exponential functions are often used to represent concentration-time profiles under the assumption of first-order kinetics, simplifying the modeling of plasma drug concentration in the body.
Linear Pharmacokinetics
Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear [Rosenbaum, 2017]. In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug. If you double the dose of a drug, the area under the curve will also double. Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses. The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics [Rosenbaum, 2017].
Figure 11: When doses of insulin are administered at close intervals, the subsequent dose may start to act before the previous dose has fully finished its effect. This can lead to a stacking of insulin effects. The figure depicts additive effects under linear pharmacokinetic assumptions.
Appendix C Additional Results
C.1 Patient-specific (Local) Models
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 4 and Fig. 5. Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig. 12.
Table 4: MAE computed across forecast horizon values using local and global models.
ID Exogenous PK
Local Global Local Global
540 12.106 10.119 11.880 10.015
(0.779) (0.139) (0.596) (0.147)
544 7.952 7.188 8.067 6.967
(0.679) (0.062) (0.490) (0.083)
552 8.559 7.400 9.132 7.374
(0.243) (0.118) (0.403) (0.070)
559 10.845 9.013 13.051 8.967
(1.266) (0.176) (7.027) (0.102)
563 9.527 8.731 9.427 8.506
(0.379) (0.065) (0.402) (0.084)
567 12.599 9.640 12.132 9.562
(1.649) (0.138) (1.667) (0.189)
570 8.361 7.523 8.466 7.258
(0.496) (0.141) (0.223) (0.121)
575 10.891 9.712 11.095 9.496
(0.419) (0.099) (0.470) (0.151)
584 11.111 9.846 11.518 9.905
(0.547) (0.091) (0.538) (0.156)
588 10.172 8.172 9.376 8.137
(1.737) (0.081) (0.321) (0.087)
591 11.987 10.246 11.428 10.151
(1.274) (0.083) (0.172) (0.134)
596 8.990 8.133 9.147 8.059
(0.246) (0.103) (0.502) (0.141)
All 10.324 8.873 10.425 8.758
(0.401) (0.064) (0.659) (0.080)
Figure 12: The NHITS-PK hybrid global-local model (Hybrid) trained on all OhioT1DM subject data (blue, solid) has a lower MAE, on average, than NHITS-PK models trained on individual subjects, or local models (purple, dashed) .
C.2 Forecast Comparison
Figure 13: Blood glucose forecasts of the final horizon point during 5 randomly selected hypoglycemic threshold periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients. The NHITS-PK hybrid global-local model (orange, dashed) produces improved forecasts of the ground truth (blue, solid) compared to the NHITS baseline model with sparse exogenous features (yellow, dash-dotted).
C.3 Forecasting Response to Treatment
Predicting treatment effects using accurate forecasting algorithms can assist individuals with diabetes and their care providers in making informed decisions about the ideal medication type and dosage, leading to improved glycemic control and reduced risk of both short- and long-term complications. Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts. We examine whether the NHITS-PK model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 14. The NHITS-PK model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
Figure 14: Treatment effect of insulin on the NHITS-PK hybrid global-local model. Each point in the plot corresponds to the final horizon point. Forecasts with original bolus doses (orange, solid), removing bolus insulin doses results in higher forecasts (blue, dashed), and augmenting doses results in lower blood glucose forecasts (yellow, dotted).
C.4 Forecasting with Dose-Independent Parameters
[…]
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters [Xie, 2018, Visentin et al., 2016]. In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
Simulated Dataset
Model All Values Critical Values
Baseline Exog. Baseline Exog.
NHITS 8.268 7.304 11.114 9.094
(0.039) (0.055) (0.237) (0.236)
NHITS-ST 7.055 8.648
(0.038) (0.068)
NHITS-PK 7.103 8.449
linear (0.069) (0.139)
NHITS-PK 7.037 8.492
nonlinear (0.093) (0.220)
OhioT1DM Dataset
NHITS 9.060 8.873 10.337 10.137
(0.142) (0.064) (0.137) (0.108)
NHITS-ST 8.958 10.131
(0.087 ) (0.100)
NHITS-PK 8.830 10.026
linear (0.060) (0.064)
NHITS-PK 8.758 9.965
nonlinear (0.080) (0.089)
C.5 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Table 6 provides complete results for the simulated data and Table 7 provides complete results for the OhioT1DM dataset.
Table 6: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exog. Baseline Exog.
MAE
ETS 9.450 13.457
(0.00) (0.00)
MLP 10.589 8.433 14.699 10.600
(2.085) (0.490) (3.180) (0.464)
RNN 13.703 20.190 21.433 35.955
(2.645) (9.966) (6.123) (23.235)
LSTM 13.557 10.321 20.537 13.550
(9.112) (2.074) (18.191) (2.817)
TCN 12.096 10.109 17.709 12.660
(1.681) (0.538) (2.995) (0.826)
NHITS 8.268 7.304 11.114 9.094
(0.039) (0.055) (0.237) (0.236)
NHITS-ST 7.055 8.648
(0.038) (0.068)
NHITS-PK 7.037 8.492
(0.093) (0.220)
RMSE
ETS 14.751 21.454
(0.00) (0.00)
MLP 15.561 11.461 22.380 14.449
(2.318) (0.553) (3.568) (0.502)
RNN 19.033 27.116 28.834 43.256
(3.479) (12.484) (6.974) (23.190)
LSTM 18.567 13.723 26.863 18.028
(11.191) (2.373) (18.642) (3.661)
TCN 17.094 13.551 25.163 16.908
(1.719) (0.628) (3.092) (1.074)
NHITS 12.918 10.546 18.436 13.371
(0.064) (0.106) (0.369) (0.387)
NHITS-ST 10.082 12.679
(0.054) (0.135)
NHITS-PK 10.039 12.339
(0.188) (0.435)
Table 7: Mean absolute error (MAE) (top) and Root Mean Square Error (RMSE) (bottom) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exog. Baseline Exog.
MAE
ETS 9.022 10.028
(0.00) (0.00)
MLP 11.065 11.120 12.630 12.698
(1.461) (0.381) (1.727) (0.821)
RNN 11.325 11.795 13.627 13.552
(1.286) (1.985) (1.573) (2.480)
LSTM 9.811 9.467 11.252 11.178
(0.570) (0.230) (0.512) (0.396)
TCN 11.525 10.760 13.647 12.678
(2.008) (0.925) (3.349) (1.459)
NHITS 9.060 8.873 10.337 10.137
(0.142) (0.064) (0.137) (0.108)
NHITS-ST 8.958 10.131
(0.087 ) (0.100)
NHITS-PK 8.758 9.965
(0.080) (0.095)
RMSE
ETS 14.569 16.225
(0.00) (0.00)
MLP 16.534 16.948 18.651 19.289
(1.280) (0.481) (1.403) (0.943)
RNN 16.498 17.066 19.446 19.423
(1.252) (2.037) (1.709) (2.532)
LSTM 14.915 14.597 16.894 16.851
(0.494) (0.279) (0.490) (0.394)
TCN 16.690 15.971 19.213 18.418
(1.891) (0.710) (2.706) (1.295)
NHITS 14.844 14.505 16.789 16.453
(0.192) (0.096) (0.189) (0.132)
NHITS-ST 14.640 16.492
(0.205) (0.203)
NHITS-PK 14.385 16.251
(0.092) (0.089)