### v8

# Global Deep Forecasting with Patient-Specific Pharmacokinetics

Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, it can be challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties of each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains in a blood glucose forecasting task using both realistically simulated and real-world data. Our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code
1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the drug dose can be observed and recorded, the body’s interaction with the drug cannot be directly observed and is typically inferred from its plasma concentration, as illustrated in Fig. 1. Consequently, the recorded drug dose alone cannot capture time-dependent drug kinetics, including processes such as absorption and elimination.
Figure 1: Insulin dose is recorded, but plasma insulin concentration is not directly measurable and typically requires invasive procedures and laboratory tests.
Consider the task of blood glucose monitoring for patients with diabetes who must regularly monitor and regulate their blood glucose level. Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospitalizations suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events. However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories. To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetic characteristics to improve the predictability of blood glucose responses to treatment.
The main contributions of this paper are:
  1. (i)
    Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
  2. (ii)
    Pharmacokinetic (PK) Encoder. We introduce a novel deep learning model-agnostic module that generates plasma concentration profiles to capture time-dependent treatment effects.
  3. (iii)
    Multiple Dose Effects Framework. We provide a computationally efficient framework for encoding the effects of multiple doses. We verify this approach mathematically under the assumption of linear pharmacokinetics.
  4. (iv)
    State-of-the-Art Results. We obtain significant accuracy improvements on large-scale simulated and real-world patient blood glucose forecasting datasets.
1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can reveal the effects of the medication over time through knowledge of drug movement within the body. For example, PK provides a well-established understanding of the changes in plasma insulin concentration over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The extent of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
2 Related Work
Blood Glucose Forecasting. Neural Forecasting has previously been used to predict blood glucose levels, outperforming traditional statistical methods. RNNs and LSTMs have been used both as models and as baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been used in proposed models and as baselines.
Modeling Blood Glucose Treatment Effects.
Prior works have looked at modeling treatment effects on blood glucose levels. Miller et al. introduced one of the first hybrid statistical-physiological models, the Deep T1D Simulator (DTD-Sim), which integrates glucose-insulin dynamics with deep learning through a deep state-space model (Miller et al., 2020). Odnoblyudova et al. improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats (Odnoblyudova et al., 2023). Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task (Odnoblyudova et al., 2023). Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
Despite advances in treatment effect modeling, current approaches face limitations, including (1) reliance on individualized, subject-specific models and (2) inflexible architectures that hinder adaptability to various base models, forecasting tasks, and model inputs. The next paragraph outlines these limitations and how our proposed contributions address them.
Previous work has developed individualized networks that overlook the benefits of developing global models that learn from various, diverse data series. Such cross-learning on diverse datasets can enhance generalization abilities, as demonstrated with Time Series Foundation Models (TSFMs) (Goswami et al., 2024). Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts. Regarding architectures, models such as the UVA / Padova simulator depend on predefined ODEs for glucose dynamics, restricting adaptability to other targets and exogenous inputs (Visentin et al., 2016; Miller et al., 2020). Similarly, deep learning approaches for sparse exogenous variables often rely on fixed base models like LSTMs (Rubin-Falcone et al., 2022). Our proposed contributions support diverse deep learning models, and are flexible to forecasting targets and input variables, ensuring future adaptability as more advanced tools emerge. Other prior work that proposes the integration of PK knowledge with deep learning focuses on domains such as long-term disease progression in cancer and COVID-19, where such PK models are not directly applicable to blood glucose and insulin kinetics. (Hussain et al., 2021; Qian et al., 2021). Our study provides a formal framework for integrating PK in deep learning under linear PK assumptions that are applicable to many drugs used in clinical practice (Rosenbaum, 2017).
3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints. Moreover, the duration of rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (Clinic, 2018; Association, n.d.). Thus, a 10-hour timespan fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations.
3.1 PK Encoder
Medication administrations are represented as instantaneous sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate specific to the patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
Figure 3: Medication administrations are represented as sparse variables in time series data. We propose a pharmacokinetic (PK) encoder to effectively capture time-dependent plasma drug concentration (Conc.).
Under linear PK assumptions, the PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.4. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude, but does not alter the PK parameters or shape of the curve.
Let $N$ be the number of subjects in the data. The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training. The initial k values are tuned as model hyperparameters.
The encoder is trained end-to-end with the model. During training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder then replaces sparse medication features in the training window with the corresponding concentration curves. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded up to and including time $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t+1:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3 and Fig. 14. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.4. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
The PK encoder training algorithm is described in Appendix A.5, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters. Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$. PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs. During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
3.3 Modeling Dose-Dependent Effects
A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017). A characteristic of this assumption is that PK parameters are constant and do not change with dose. In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site. Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose. For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018). This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses. However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs. 12 and 13 in Appendix A.1. To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity, so that we can capture potential PK variations without resorting to a fully nonlinear PK model. This method simplifies the modeling process by maintaining linear assumptions within each dose category. It avoids the complexity while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics. Specifically, we separately estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values. For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent kinetics. However, all physiological models approximate complex biological processes. Our goal is to develop models validated under established linear PK assumptions from the pharmacology literature while still accounting for dose-dependent effects, balancing simplicity and precision. This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
The equations presented are shown in a patient-specific manner to highlight that the PK encoder generates individualized information. However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.
3.4 Models
To assess forecasting performance across various model architectures, we trained global models with 11 different algorithms as baselines. Additionally, to assess our proposed hybrid global-local architecture, we compare the best performing model with the PK encoder to both purely local and global models.
Statistical and Deep Learning Baselines.
We employed Exponential Smoothing (AutoETS; Hyndman 2008) as our statistical baseline model. Our deep learning baselines include models from various architectural categories, such as Linear models, Multi-Layer Perceptron (MLP)-based, Recurrent Neural Network (RNN)-based, Convolutional Neural Network (CNN)-based, and Transformer-based models. We employ the following deep learning models: DLinear (Zeng et al., 2023), MLP (Rosenblatt, 1958), NHITS (Challu et al., 2022), NBEATSx (Oreshkin et al., 2020; Olivares et al., 2022a), RNN (Elman, 1990; Cho et al., 2014), Long-Short-Term Memory (LSTM; Sak et al. (2014)), Temporal Convolutional Network with an MLP decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)), Informer (Zhou et al., 2021), patch time series Transformer (PatchTST; Nie et al. (2023)), and Temporal Fusion Transformer (TFT; Lim et al. (2021)). More information on these models and related prior work is provided in Appendix A.2. Details on model training and hyperparameter selection are provided in Appendix A.3.
Global Model Baselines.
Global models are trained on all patient data to create a single model. Global model baselines include: (1) univariate models trained with only the blood glucose signal and static features, (2) models trained with univariate blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with the univariate blood glucose signal, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022). Univariate models are included as baselines to highlight forecast improvements from training on historical exogenous variables, motivating advanced transformations for sparse, continuous exogenous features using methods like our PK encoder.
With the exception of TFT, most Transformer-based models, including Informer and PatchTST, are not designed to model covariates. For these two models, we can only report global model results for those trained on univariate blood glucose data. Since our proposed PK encoder is intended to supplement model architectures capable of incorporating exogenous variables alongside the endogenous signal, we focus on evaluating the PK encoder in top-performing models that are capable of modeling covariates, such as TFT, NBEATSx, and NHITS. Beyond modeling covariates, residualized MLP-based models like NBEATSx and NHITS are valuable to demonstrate the utility of our PK encoder as NHITS has achieved SOTA performance, outperforming RNNs and Transformers in long-range forecasting tasks while avoiding the quadratic complexity of self-attention in Transformer-based architectures (Challu et al., 2022; Ansari et al., 2024).
Local Model Baselines.
Patient-specific sparse exogenous models as well as patient-specific exogenous models using the PK encoder were compared against our hybrid global-local PK encoder model.
3.5 Model Training
Models were trained using Adam optimizer (Kingma and Ba, 2017). Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)=\left\{\begin{array}[]{ll}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{array}\right.$ (8)
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. Optimal hyperparameters were selected based on a validation set with cross-validation. More information on model training and hyperparameters in provided in Appendix A.3.
3.6 Model Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE):
$\mathrm{MAE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\frac{1}{H}\sum_{{\tau}=t+1}^{t+H}|\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau}|\quad\text{and}$ (9)
$\mathrm{RMSE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\sqrt{\frac{1}{H}\sum_{\tau=t+1}^{t+H}(\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau})^{2}}.$ (10)
We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models. We also evaluate the performance of the models exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation. To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-test statistics computed on the patients’ mean MAE results.
4 Cohort
We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements. These open-source datasets ensure reproducibility and are widely used in blood glucose forecasting research, featuring some of the largest patient cohorts (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
Simulated data.
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. Appendix A.1 contains examples of blood glucose time series for 6 individuals, shown in Fig. 10, as well as basal and bolus dose distributions for 12 individuals, shown in Fig. 12.
Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower average MAE than NHITS-PK local models (blue, dashed median).
Figure 6: The NHITS-PK hybrid global-local model (orange, solid median) consistently outperforms the local NHITS-PK model (blue, dashed median).
[…]
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 9.450 13.457 9.022 10.028
DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
Informer 8.599 12.074 13.714 14.867
PatchTST 7.284 9.842 9.922 10.999
TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
NHITS 8.268 7.304 11.114 9.094 9.060 8.873 10.337 10.137
TFT-PK 5.743 7.042 9.529 10.461
NBEATSx-PK 6.806 7.998 9.983 10.927
NHITS-PK 7.037 8.492 8.758 9.965
Real-world data.
Twelve de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. Missing values were forward filled to prevent data leakage. The sample count and percentages of missing data for each subject are included in Table 4 in Appendix A.1. Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig. 11, as well as basal and bolus dose distributions for the 12 patients, shown in Fig. 13.
5 Results
PK-encoder models outperform their sparse exogenous model counterparts.
For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values. The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual subjects. For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values. The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients. Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.3 Tables 8 and 9, respectively.
Hybrid global-local models significantly improve forecasts over global models.
For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for subjects for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts. For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for subjects for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts. Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
Hybrid global-local models significantly improve forecasts over local models.
For the OhioT1DM dataset, the NHITS-PK hybrid global-local model outperforms local NHITS-PK models developed for individual OhioT1DM patients by 15.8%, on average (p-value: 1.7e-5) as shown in Fig. 5. Furthermore, this improvement in forecast accuracy is consistent across all patients as shown in Fig. 6 and Table 3 (standard deviations are provided in Table 6 in Appendix C.1). Our hybrid global-local NHITS-PK model also significantly outperforms local sparse exogenous baseline models by 14.8%, on average, for individual subjects (p-value: 1.7e-6).
Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. Best result is in bold.
Simulated Dataset OhioT1DM Dataset
TFT-PK TFT-ST NHITS-PK NHITS-ST
All Values 5.743 5.885 8.758 8.958
(0.159) (0.187) (0.080) ((0.087)
Critical Values 7.042 7.233 9.965 10.131
(0.270) (0.282) (0.095) (0.100)
Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. Best results in bold.
ID w/ Exog. (Sparse) w/ Exog. (PK)
Local Global Local Hybrid
540 12.106 10.119 11.880 10.015
544 7.952 7.188 8.067 6.967
552 8.559 7.400 9.132 7.374
559 10.845 9.013 13.051 8.967
563 9.527 8.731 9.427 8.506
567 12.599 9.640 12.132 9.562
570 8.361 7.523 8.466 7.258
575 10.891 9.712 11.095 9.496
584 11.111 9.846 11.518 9.905
588 10.172 8.172 9.376 8.137
591 11.987 10.246 11.428 10.151
596 8.990 8.133 9.147 8.059
All 10.324 8.873 10.425 8.758
The NHITS-PK model effectively learns the inhibitory treatment effects of insulin on blood glucose.
We examine whether the NHITS-PK model can accurately capture treatment effects of insulin on blood glucose by generating forecasts for three scenarios: the original bolus insulin doses, bolus insulin doses removed, and bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 8. This result underscores the importance of models that can effectively capture treatment effects, enabling them to adapt to new treatment doses for more reliable and accurate forecasts.
Figure 7: Blood glucose forecasts of the final horizon point during 5 randomly selected hyperglycemic periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients. Forecasts for the NHITS-PK model (orange, dashed) are more aligned with ground truth (black, solid) compared to the NHITS sparse exogenous model blue, dash-dotted).
Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed). Each point represents the final forecasted timestamp from a rolling window forecast.
6 Discussion
The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data. These improvements are not only statistically significant, but they have potential for positive real-world clinical impact. In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig. 9. This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations. Improving the prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009). Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing. While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge. It can also be easily adapted for other types of medication administration, such as oral or intravenous injections.
In addition to improving performance over global model counterparts, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models. This result is of substantial importance in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts. Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models. Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space. The hybrid model also has the potential to be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
Our study demonstrates that leveraging domain knowledge to integrate exogenous variables in models offers performance benefits over using sparse features or the “Sum Total” feature engineering approach, both of which are outperformed by our proposed deep PK encoding method. Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
Limitations and Future Work.
This study has two limitations related to the PK encoder: we assume that the concentration curves exhibit 100% bioavailability and maintain a consistent peak concentration across patients. These simplifying assumptions are intended to focus model learning on the parameter $k$, rather than on the area or peak of the concentration curve, to emphasize the spread or duration of the concentration curve and to reduce potential issues with parameter identification. Future work will explore expanding concentration curve parameters to estimate area and peak parameters to potentially enable more flexible concentration curves. We will also aim to apply our approach to additional cohorts of patients, such as those with T2DM, as more data become available, to further validate our findings. Future work will also explore modeling dose-dependent effects using more dynamic nonlinear PK frameworks.
Figure 9: The PK encoder enables improved NHITS model prediction lead times for hyperglycemic events (CGM $\geq$ 180) for 9 of 12 OhioT1DM patients.
[…]
This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
Data and Code Availability
We use open-source simulated and real-world data that is available to other researchers. The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository (Xie, 2018; Visentin et al., 2016). The real-world data includes the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu (2020). Implementations of models used in our work were obtained from the open-source Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries. We implemented our method by adopting the Neuralforecast repository to leverage their model training framework. Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
Appendix A Supplemental Methods
A.1 Data
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
The ‘Dexcom’ continuous glucose monitor (CGM) option was used to obtain glucose readings at 5-minute intervals. The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio. However, we adapt the controller to deliver basal doses once per hour, consistent with the function of real insulin pumps used in the OhioT1DM dataset. The resulting simulated data consists of CGM (in mg/dL) values, insulin (units), and carbohydrates (in grams), referred to as CHO. We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses. Blood glucose time series for 6 simulated patients are shown in Fig. 10 in Appendix A.1. Basal and bolus dose distributions for 12 simuluted patients are shown in Fig. 12 in Appendix A.1.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
Sparse exogenous features in the data include basal rate (in units per hour), temporary basal rate (units per hour), and bolus insulin (units), and CHO values. Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period. ‘Square dual’ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps. If multiple bolus administrations occurred within the same minute, the insulin units were summed. For Medtronic insulin pumps 530G and 630G versions, basal insulin is administered hourly based on the basal rate. For timestamps with overlapping basal rates, the last basal rate value was considered. When specified, the temporary basal rate superseded the basal rate.
We use the specified train and test partitions provided in the dataset. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. The sample count and percent missing data for each subject are included in Table 4 in Appendix A.1. Missing values were forward filled to prevent data leakage. Blood glucose time series plots for 6 subjects are shown in Fig. 11 in Appendix A.1. Basal and bolus dose distributions for the 12 patients are shown in Fig. 13 in Appendix A.1.
Fig. 10 shows the first 10 days of data for 6 subjects from the simulated dataset. Fig. 11 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset. Figs. 12 and 13 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively. Table 4 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 4: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
ID Original Preprocessed
Train Test Test
540 13109 3066 2691
(8.87%) (6.43%) (6.91%)
544 12671 3137 2691
(16.16%) (13.42%) (14.08%)
552 11097 3950 2691
(18.80%) (40.15%) (53.36%)
559 12081 2876 2691
(10.64%) (12.59%) (13.45%)
563 13098 2691 2691
(7.44%) (4.50%) (4.50%)
567 13536 2871 2691
(19.78%) (16.79%) (16.91%)
570 11611 2880 2691
(5.42%) (4.69%) (4.91%)
575 13103 2719 269
(9.44%) (4.74%) (4.79%)
584 13248 2995 2691
(8.29%) (11.02%) (11.93%)
588 13105 2881 2691
(3.55%) (3.12%) (3.34%)
591 12755 2847 2691
(14.96%) (3.06%) (3.23%)
596 13630 3003 2691
(20.20%) (8.66%) (9.66%)
A.2 Models
The baseline models used in this work have been extensively used in previous blood glucose forecasting work. For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020). Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020) have been used in proposed models and as baselines.
Exponential Smoothing (AutoETS) The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
DLinear (DLinear) A Linear model that leverages linear layers to model the trend and seasonal components. A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
Neural Basis Expansion Analysis with Exogenous (NBEATSx) An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections. The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
Informer (Informer) (Zhou et al., 2021) A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network. A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity.
Patch time series Transformer (PatchTST) (Nie et al., 2023) A Transformer-based deep learning model with a multi-head attention mechanism. Input time series into sub-series level patches as input tokens.
Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
A.3 Model Training and Hyperparameters
The deep learning models were trained using an NVIDIA A100 Tensor Core GPU. For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 5. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 5. We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
Hyperparameter Considered Values
Learning rate Loguniform(1e-4, 1e-1)
Batch size 4
Windows batch size’ 256
Dropout 0.0
Training steps 2000
Validation check steps 100
Early stop 5
patience steps
Random seed DiscreteRange(1, 10)
Initial $\textbf{k}_{\text{basal}}$1.0, 1.1, 1.2}
Initial $\textbf{k}_{\text{bolus}}$1.7, 1.8, 1.9}
Initial $\textbf{k}_{\text{CHO}}$1.7, 1.8, 1.9}
Table 5: Common hyperparameter search space
We tune additional specific hyperparameters for each architecture. For the MLP model, we tune the number of hidden units per layer ([128, 256]). For the LSTM model, we tune the hidden size ([128, 256]). For the TCN model, we tune the hidden size ([128, 256]) and number of dilations ([2, 4, 8]). For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]). For the Informer model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and number of decoder layers ([1, 2]). For the TFT model, we tune the hidden size ([128, 256, 512]) and number of attention heads ([4, 8, 16]). For the NBEATSx and NHITS models, we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Propositions
Proposition A.1.
Under linear pharmacokinetic (PK) assumptions, the area under the concentration curve is directly proportional to the dose of the drug. As such, PK parameters, such as bioavailability, are assumed to be constant for a given drug. The bioavailability, $F$, can be assessed from the dose-normalized area under the concentration curve (AUC). When bioavailability is constant, we show that the AUC is bounded by the dose, $x$, given that $0\leq F\leq 1$,
$F$ $\propto\frac{\text{AUC}}{x}$ (11)
$n\cdot F$ $=n\cdot\frac{\text{AUC}}{x}$ (12)
AUC $=F\cdot x$ (13)
$\leq x$ (14)
where $n$ where is a constant of proportionality.
Proposition A.2.
Given some mathematical model, $C(t,\cdot)_{x}$, of concentration for a drug dose, $x$, that is a function of time, $t$, and pharmacokinetic parameters, $\cdot$, we show that the overall effect (concentration) of $n$ drug doses of the same drug is equal to the sum of the individual effects under linear pharmacokinetic assumptions:
[…]
$=\int_{0}^{T}\sum_{i=1}^{n}C(t,\cdot)_{x}dt$ (19)
Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
A.5 Algorithm
We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles. This framework can be implemented in any deep learning architecture. It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch. This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018). For each training step, the algorithm iterates over batches of data. It computes a context vector c based on the input data x, k, and other parameters. It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values. After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step. It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$. It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss. Finally, the algorithm returns the final k values.
Algorithm 1 PK Encoder Training.
$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
Result: k
$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
for each step i from 1 to E do
       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
for batch $b\in\mathcal{B}_{train}$ do
               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
       end for
      if i % validation step == 0 then
               $\ell(\textbf{k})\leftarrow 0$
for batch $b\in\mathcal{B}_{val}$ do
                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
             end for
            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
       end if
end for
return k
Figure 10: Time series for the first 10 days of 6 subjects from the simulated dataset.
Figure 11: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
Figure 12: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
Figure 13: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
Appendix B Pharmacokinetic Background
Concentration-time Profiles
From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects (Rosenbaum, 2017; Angelika et al., 2005). Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile. Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves. The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig. 2, along with various patient-specific factors.
Insulin Analogs
Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action. Fig. 2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs (Angelika et al., 2005). Generally, bolus doses are quick-acting and, thus, often taken before meals. Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day. However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin. For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes. Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
Bioavailability
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012). However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
First-order Kinetics
First-order kinetics, a common assumption in pharmacokinetics, describes the rate at which a drug is eliminated from the body as being proportional to its concentration. This means that a constant fraction of the drug is eliminated per unit of time, regardless of its concentration levels. Exponential functions are often used to represent concentration-time profiles under the assumption of first-order kinetics, simplifying the modeling of plasma drug concentration in the body.
Linear Pharmacokinetics
Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear (Rosenbaum, 2017). In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug. If you double the dose of a drug, the area under the curve will also double. Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses. The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics (Rosenbaum, 2017).
Figure 14: When doses of insulin are administered at close intervals, the subsequent dose may start to act before the previous dose has fully finished its effect. This can lead to a stacking of insulin effects. The figure depicts additive effects under linear pharmacokinetic assumptions.
Appendix C Additional Results
C.1 Patient-specific (Local) Models
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 6 and Fig. 6. Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig. 5.
Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. Best results are in bold.
ID w/ Exog. (Sparse) w/ Exog. (PK)
Local Global Local Global
540 12.106 10.119 11.880 10.015
(0.779) (0.139) (0.596) (0.147)
544 7.952 7.188 8.067 6.967
(0.679) (0.062) (0.490) (0.083)
552 8.559 7.400 9.132 7.374
(0.243) (0.118) (0.403) (0.070)
559 10.845 9.013 13.051 8.967
(1.266) (0.176) (7.027) (0.102)
563 9.527 8.731 9.427 8.506
(0.379) (0.065) (0.402) (0.084)
567 12.599 9.640 12.132 9.562
(1.649) (0.138) (1.667) (0.189)
570 8.361 7.523 8.466 7.258
(0.496) (0.141) (0.223) (0.121)
575 10.891 9.712 11.095 9.496
(0.419) (0.099) (0.470) (0.151)
584 11.111 9.846 11.518 9.905
(0.547) (0.091) (0.538) (0.156)
588 10.172 8.172 9.376 8.137
(1.737) (0.081) (0.321) (0.087)
591 11.987 10.246 11.428 10.151
(1.274) (0.083) (0.172) (0.134)
596 8.990 8.133 9.147 8.059
(0.246) (0.103) (0.502) (0.141)
All 10.324 8.873 10.425 8.758
(0.401) (0.064) (0.659) (0.080)
C.2 Forecasting with Dose-Independent Parameters
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180). Result for models with (w/) and without (w/o) exogenous variables are shown. Best results are in bold.
Simulated Dataset
Model All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog.
NHITS 8.268 7.304 11.114 9.094
(0.039) (0.055) (0.237) (0.236)
NHITS-ST 7.055 8.648
(0.038) (0.068)
NHITS-PK 7.103 8.449
linear (0.069) (0.139)
NHITS-PK 7.037 8.492
nonlinear (0.093) (0.220)
OhioT1DM Dataset
NHITS 9.060 8.873 10.337 10.137
(0.142) (0.064) (0.137) (0.108)
NHITS-ST 8.958 10.131
(0.087 ) (0.100)
NHITS-PK 8.830 10.026
linear (0.060) (0.064)
NHITS-PK 8.758 9.965
nonlinear (0.080) (0.089)
[…]
We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values. However, this approach led to worse performance on the OhioT1DM dataset. The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters (Xie, 2018; Visentin et al., 2016). In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3. Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data. These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
C.3 Model MAE and RMSE Results
This appendix presents the complete results for all models for both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis. Tables 8 and 9 provide complete MAE and RMSE results, respectively.
Table 8: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset. Average results are computed across 8 trials. Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based, and PK model groups, respectively. Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable. Best results are highlighted in bold, second best results are underlined. PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 9.450 13.457 9.022 10.028
(0.00) (0.00) (0.00) (0.00)
DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
(0.054) (0.141) (0.126) (0.169) (1.595) (3.590) (1.228) (4.080)
RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
(2.645) (9.966) (6.123) (23.235) (1.286) (1.985) (1.573) (2.480)
LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
(9.112) (2.074) (18.191) (2.817) (0.570) (0.230) (0.512) (0.396)
TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
(1.681) (0.538) (2.995) (0.826) (2.008) (0.925) (3.349) (1.459)
Informer 8.599 12.074 13.714 14.867
(1.060) (2.083) (3.753) (3.620)
PatchTST 7.284 9.842 9.922 10.999
(0.554) (0.939) (0.949) (1.054)
TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
(0.115) (0.192) (0.178) (0.243) (0.811) (0.455) (1.038) (0.792)
MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
(2.085) (0.490) (3.180) (0.464) (1.461) (0.381) (1.727) (0.821)
NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
(0.236) (0.031) (0.547) (0.072) (0.528) (0.214) (0.548) (0.188)
NHITS 8.268 7.304 11.114 9.094 9.060 8.873 10.337 10.137
(0.039) (0.055) (0.237) (0.236) (0.142) (0.064) (0.137) (0.108)
TFT-PK 5.743 7.042 9.529 10.461
(0.159) (0.270) (0.597) (0.597)
NBEATSx-PK 6.806 7.998 9.983 10.927
(0.060) (0.076) (0.264) (0.271)
NHITS-PK 7.037 8.492 8.758 9.965
(0.093) (0.220) (0.080) (0.095)
Table 9: Root Mean Square Error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset. Average results are computed across 8 trials. Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based groups, and PK model groups, respectively. Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable. Best results are highlighted in bold, second best results are underlined. PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
Simulated Dataset OhioT1DM Dataset
Model All Values Critical Values All Values Critical Values
w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
AutoETS 14.751 21.454 14.569 16.225
(0.00) (0.00) (0.00) (0.00)
DLinear 12.545 12.673 17.606 17.707 17.138 18.609 18.679 20.743
(0.085) (0.119) (0.340) (0.216) (1.462) (3.511) (0.970) (3.869)
RNN 19.033 27.116 28.834 43.256 16.498 17.066 19.446 19.423
(3.479) (12.484) (6.974) (23.190) (1.252) (2.037) (1.709) (2.532)
LSTM 18.567 13.723 26.863 18.028 14.915 14.597 16.894 16.851
(11.191) (2.373) (18.642) (3.661) (0.494) (0.279) (0.490) (0.394)
TCN 17.094 13.551 25.163 16.908 16.690 15.971 19.213 18.418
(1.719) (0.628) (3.092) (1.074) (1.891) (0.710) (2.706) (1.295)
Informer 12.348 17.400 18.536 19.904
(0.918) (1.865) (3.459) (3.220)
PatchTST 11.467 15.755 15.375 17.065
(0.603) (1.000) (0.838) (0.928)
TFT 10.691 8.898 14.188 10.979 14.832 14.571 16.523 16.577
(0.139) (0.314) (0.329) (0.446) (0.791) (0.345) (0.959) (0.684)
MLP 15.561 11.461 22.380 14.449 16.534 16.948 18.651 19.289
(2.318) (0.553) (3.568) (0.502) (1.280) (0.481) (1.403) (0.943)
NBEATSx 14.105 9.570 20.526 11.397 15.934 16.224 17.540 17.742
(0.348) (0.053) (0.756) (0.104) (0.548) (0.675) (0.599) (0.303)
NHITS 12.918 10.546 18.436 13.371 14.844 14.505 16.789 16.453
(0.064) (0.106) (0.369) (0.387) (0.192) (0.096) (0.189) (0.132)
TFT-PK 8.620 10.517 14.725 16.265
(0.188) (0.367) (0.414) (0.437)
NBEATSx-PK 9.480 11.244 16.945 17.659
(0.084) (0.117) (1.467) (0.658)
NHITS-PK 10.039 12.339 14.385 16.251
(0.188) (0.435) (0.092) (0.089)